OpenStudy (anonymous):

You have a job that pays $ 6.20 an hour. You get $ 8.30 an hour for overtime (anything over 40 hours). How many hours do you need to work to earn $289.5? Answer in units of hours

OpenStudy (bahrom7893):

6.20x+8.30(40+x)=289.5

OpenStudy (bahrom7893):

Solve for x

OpenStudy (bahrom7893):

hmm wait that's wrong..

OpenStudy (anonymous):

40*6.20=248, so you need to work longer than 40 hours

OpenStudy (anonymous):

289.5-248=41.5 41.5/8.3= 5 So you need to work a total of 40+5=45 hours.